OpenStudy (anonymous):

Kyle drives two and three-fourths hours to visit his grandmother when he takes the new expressway, which has cut his trip time in half. How long did his trip take before the expressway was built? five and one-half hours one and one-third hours five-sixths hours two and one-third hours

OpenStudy (anonymous):

@evilmath

OpenStudy (anonymous):

two and three-fourths plus two and three-fourths.

OpenStudy (anonymous):

\[2\frac{ 3 }{ 4 } = \frac{ 11 }{ 4 }; \frac{ 11 }{ 4 }+ \frac{ 11 }{ 4 } = \frac{ 22 }{ 4 } = 5\frac{ 2 }{ 4 } = 5 and one half\]

OpenStudy (anonymous):

@evilmath the girl who answered the questions on yahoo only did it for a couple of problems not this one..

OpenStudy (anonymous):

@DoYourHomework ty